# Robust Plan Evaluation based on Approximate Probabilistic Machine Learning

### v2

@@ 6. Related Work
Re-optimization Techniques use intervals instead of point estimates to identify robust plans within specific parameter ranges (Babu et al., 2005; Markl et al., 2004; Moumen et al., 2016). These plans are used at run-time while actual parameter values or precise intervals are discovered either via execution or sampling. When these values are outside of the robustness range of the plan, the optimizer suggests another plan expected to be robust in this range. The execution of the previous plan terminates prematurely and switches to the new plan. These approaches necessitate intrusive modifications to extend the execution plans and the execution engine to allow for re-optimization.
@@
Discovery-based Techniques eliminate reliance on parameter estimations by exploring the entire parameter space during compilation. They use mechanisms to iteratively execute plans within time budgets, discarding results when a plan exceeds its allocated time (Dutt and Haritsa, 2016; Karthik et al., 2016, 2018). These methods suffer from significant compilation overheads, making them suitable primarily for OLAP-style environments. Also, they assume perfect correlation between the optimizer’s cost function and execution time, and they require partial execution of suboptimal plans to discover parameter values.
@@
Robustness Quantification Techniques incorporate robustness alongside optimality in plan selection. One work uses ”Least Expected Cost” instead of ”Least Specific Cost,” deriving a PDF of plan costs (Chu et al., 2002). This enables selection based on the expected cost across the parameter space rather than a specific parameter setting. Deriving the distribution of cost analytically, however, is not a trivial task given the large number of error-prone parameters involved. Other works propose a framework to estimate an upper bound for cost given cardinality estimation errors (Kastrati and Moerkotte, 2016; Moerkotte et al., 2009). While this can quantify plan robustness, it depends on prior knowledge of the cardinality errors. Another work evaluates robustness based on the derivative and area under the curve of the PCF (Wolf et al., 2018). While this approach is objective, it overlooks the probabilistic nature of estimates and assumes uniform distributions of parameters. (Hertzschuch et al., 2021). In addition, it relies on classical cost models built by making simplifying assumptions (Christodoulakis, 1984), potentially leading to poor performance even with accurate parameter estimates. Moreover, this approach assumes the PCF is a linear function which does not hold in practice.
@@
Adaptive Techniques delay some optimization decisions to runtime when true statistics are obtained. Adaptive Join Algorithm (AJA) (Sack, 2019; Zhang et al., 2023) switches a hash join to a nested-loops join after determining the true cardinality of the built hash table. If the cardinality is below a threshold it switches to a nested-loops join. Look-ahead Information Passing (LIP) (Zhu et al., 2017) generalizes a semijoin optimization to a pipeline of equijoins typically observed in left-deep plans based on star schema queries. Smooth Scan (Borovica-Gajic et al., 2018) proposes a new access operator that morphs between a sequential scan and an index scan depending on the statistics obtained at runtime. These approaches reduce sensitivity to cardinality misestimations (Zhang et al., 2023). They are complementary to learning-based techniques as they improve query processing at runtime while learning-based approaches improve query optimization.
@@
Learning-based Techniques aim to learn the execution cost from runtime (Hilprecht and Binnig, 2022; Marcus et al., 2019, 2024; Siddiqui et al., 2020). These works employ learning from a set of queries and corresponding plans using either supervised or reinforcement learning. While one study shows some robustness with respect to estimation errors (Marcus et al., 2019), these works still face uncontrolled predictive uncertainties inherent in learning-based methods. In addition, they do not address the issue of potential performance regressions compared to classic optimizers. To address this issue, Hybrid Cost Model (Wang et al., 2023) proposes a meta-model that learns to route queries to suitable cost models (either learned or classic ones) when they are expected to produce better plans, but it does not account for uncertainties or show robustness to out-of-distribution data. A similar study (Yu et al., 2022) proposes a hybrid approach that combines learning-based cost models with traditional optimizers, selecting partial left-deep join orders based on the predictions of the model. It uses uncertainty estimates to eliminate highly uncertain plans, with a fallback to the classic optimizer when uncertainty is high for all plans in the search space. This work also does not demonstrate the robustness of the produced plans or the robustness to out-of-distribution samples. Additionally, it does not acknowledge that a highly uncertain plan may still offer substantially lower costs than a robust plan and that pruning may not always be beneficial. In comparison, Roq’s risk-aware strategies use uncertainties not only for pruning but also for plan evaluation and selection by employing suboptimality risk quantification or conservative plan selection strategies. In addition, they allow one to prioritize robustness over optimality (or vice versa).Furthermore, they demonstrate strong robustness not only for unseen queries but also for out-of-distribution samples.
@@ 7. Conclusion
In this paper, we present Roq, a novel approach based on approximate probabilistic ML that enhances the robustness of query optimization in database systems. We establish a theoretical framework to formalize the concepts of robustness and risk in the context of plan evaluation and selection. Our framework includes methodologies for quantifying plan and estimation risks, as well as assessing the risks associated with evaluating and selecting a plan using approximate probabilistic ML. We propose strategies for plan evaluation and selection that leverage risk measures and introduce an innovative risk-aware learned cost model that incorporates query and plan encoding techniques, based on GNNs, to effectively capture pertinent information and associations between query and plan features. Roq is subjected to a comprehensive experimental study, demonstrating superior performance compared to the state-of-the-art techniques in terms of predictive accuracy and robustness while improving runtime performance, too.

### v3

@@ 6. Related Work
Re-optimization Techniques use intervals instead of point estimates to identify robust plans within parameter ranges (Babu et al., 2005; Markl et al., 2004; Moumen et al., 2016). These plans are used at run-time, while actual parameter values or intervals are discovered either via execution or sampling. When values are outside of the robustness range of the plan, the optimizer suggests another plan expected to be robust in the new range. The plan execution terminates prematurely and switches to the new plan. These approaches necessitate intrusive modifications to extend the plans and the execution engine to allow for re-optimization.
@@
Discovery-based Techniques eliminate reliance on parameter estimations by exploring the entire parameter space during compilation. They use mechanisms to iteratively execute plans within time budgets, discarding results when a plan exceeds its allocated time (Dutt and Haritsa, 2016; Karthik et al., 2016, 2018). These methods suffer from significant compilation overheads, making them suitable primarily for OLAP-style environments. Also, they assume a perfect correlation between the optimizer’s cost function and the execution time, and they require partial execution of suboptimal plans to discover parameter values.
@@
Robustness Quantification Techniques incorporate robustness alongside optimality in plan selection. One work uses ”Least Expected Cost” instead of ”Least Specific Cost”, deriving a PDF of plan costs (Chu et al., 2002). This enables selection based on the expected cost across the parameter space rather than a specific parameter setting. Deriving the distribution of cost analytically, however, is not a trivial task given the large number of error-prone parameters involved. Other works propose a framework to estimate an upper bound for cost given cardinality estimation errors (Kastrati and Moerkotte, 2016; Moerkotte et al., 2009). While this can quantify plan robustness, it depends on prior knowledge of the cardinality errors. Another work evaluates robustness based on the derivative and area under the curve of the PCF (Wolf et al., 2018). This approach overlooks the probabilistic nature of the estimated parameters and assumes they are uniformly distributed. (Hertzschuch et al., 2021). In addition, it relies on classical cost models built by making simplifying assumptions (Christodoulakis, 1984), potentially leading to poor performance even with accurate parameter estimates. Moreover, this approach assumes the PCF is a linear function, which does not hold in practice.
@@
Adaptive Techniques delay some optimization decisions to runtime, when true statistics are obtained. Adaptive Join Algorithm (AJA) (Sack, 2019; Zhang et al., 2023) switches a hash join to a nested-loops join after determining the true cardinality of the built hash table. Look-ahead Information Passing (LIP) (Zhu et al., 2017) generalizes a semijoin optimization to a pipeline of equijoins, typically observed in left-deep plans based on star schema queries. Smooth Scan (Borovica-Gajic et al., 2018) proposes a new access operator that morphs between a sequential and an index scan depending on statistics at runtime. These approaches reduce sensitivity to cardinality misestimations (Zhang et al., 2023). They are complementary to learning-based techniques, as they improve query processing at runtime while the latter improve query optimization.
@@
Learning-based Techniques aim to learn the execution cost from runtime (Hilprecht and Binnig, 2022; Marcus et al., 2019, 2024; Siddiqui et al., 2020). These works employ learning from a set of queries and corresponding plans using supervised or reinforcement learning. While one study shows some robustness with respect to estimation errors (Marcus et al., 2019), these works still face uncontrolled predictive uncertainties inherent in learning-based methods. They do not address the issue of potential performance regressions compared to classic optimizers. Hybrid Cost Model (Wang et al., 2023) proposes a meta-model that learns to route queries to suitable cost models (either learned or classic ones) when they are expected to produce better plans, but it does not account for uncertainties or show robustness to out-of-distribution data. A similar study (Yu et al., 2022) proposes a hybrid approach that combines learning-based cost models with traditional optimizers, selecting partial left-deep join orders based on the predictions of the model. It uses uncertainty estimates to eliminate highly uncertain plans, with a fallback to the classic optimizer when all plans are highly uncertain. It does not demonstrate the robustness of produced plans or robustness to out-of-distribution samples. Also, it does not acknowledge that a highly uncertain plan may still be substantially cheaper than a robust plan and that pruning may not be beneficial. In comparison, Roq’s risk-aware strategies use uncertainties not only for pruning but also for plan evaluation and selection by employing suboptimality risk quantification or conservative plan selection strategies. In addition, they allow one to prioritize robustness over optimality (or vice versa). Also, they demonstrate strong robustness not only for unseen queries but also for out-of-distribution samples.
@@ 7. Conclusion
We present Roq, a novel approach based on approximate probabilistic ML that enhances the robustness of query optimization in RDBMSs. We establish a theoretical framework to formalize the concepts of robustness and risk for plan evaluation and selection, including techniques for quantifying plan and estimation risks. We propose risk-aware strategies for plan evaluation and selection that leverage these risk measures, along with a learned cost model incorporating GNN-based query and plan embeddings. In comprehensive experiments, Roq demonstrates superior predictive accuracy, improved robustness to workload shifts, and significant performance gains over state-of-the-art approaches, while maintaining practical compilation overheads.